OpenStudy (anonymous):

find an equation of tangent line given a point Q(9,5) with center P(6,4)

OpenStudy (amistre64):

tangent to a point is not defined

OpenStudy (anonymous):

?

OpenStudy (amistre64):

read your post and see if you can make any sense out of it ....

OpenStudy (anonymous):

well i have the standard form which is (x-6)^2+(y-4)^2=square root of 10

OpenStudy (amistre64):

thats something we can work with then :)

OpenStudy (anonymous):

thank you

OpenStudy (amistre64):

take its derivative ...

OpenStudy (anonymous):

im not in calculus yet

OpenStudy (anonymous):

im in pre-calculus

OpenStudy (amistre64):

then we need to find the slope of the line from the Q to P which then makes sense

OpenStudy (amistre64):

the tangent will be the perp slope

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

show me

OpenStudy (anonymous):

center = (6,4)

OpenStudy (amistre64):

Q(9,5) with center P(6,4) -9-5 -9-5 ------- ------ 0,0 -3,-1 ; slope = y/x: 1/3

OpenStudy (anonymous):

point = (9,5)

OpenStudy (amistre64):

in this case; either point doesnt matter; slope between them is still the same

OpenStudy (anonymous):

why is -9 and not just 9?

OpenStudy (amistre64):

becasue 9+9 isnt 0

OpenStudy (amistre64):

to get (9,5) to (0,0) we have to subtract it from itself

OpenStudy (amistre64):

and whatever we do to one point, we do to them all

OpenStudy (anonymous):

ugh....

OpenStudy (anonymous):

you lost me there

OpenStudy (anonymous):

sorry its just that i haven't done math in so long

OpenStudy (amistre64):

thats becasue im doing this my way; how do you find slope between 2 points?

OpenStudy (anonymous):

ill medal you each answer

OpenStudy (anonymous):

change in y/change in x

OpenStudy (amistre64):

good; then what is our change in y and our change in x from point to point?

OpenStudy (anonymous):

so i would do 9-5/5-4 = 4/1 which is equal to just 4

OpenStudy (amistre64):

subtract one point from the other and you end up with the changes

OpenStudy (amistre64):

your slope formula is fine except that you are not placing the points in the right spot which is why i avoid that method and just step it out....

OpenStudy (amistre64):

Q(9,5) -P(6,4) ------ 3, 1 ; slope = 1/3

OpenStudy (anonymous):

oh sorry

OpenStudy (anonymous):

i mean 9-6/5-4 = 3/1 = 3

OpenStudy (amistre64):

your 9-5 should have been 9-6 ... but becasue of the manner in which you are trying to fill in parts of a formula you tend to bring in more confusion along the way

OpenStudy (amistre64):

and your putting your xs over your ys

OpenStudy (anonymous):

oh wait y/x

OpenStudy (anonymous):

so its 5-4/9-6 = 1/3

OpenStudy (amistre64):

subtract your points; then stack y/x ... avoids alot of messups

OpenStudy (anonymous):

ugh im so stupid

OpenStudy (amistre64):

1/3 yes is the slope between the points; now flip and negate for a perp

OpenStudy (amistre64):

-3/1 = -3

OpenStudy (anonymous):

okay now that i have the slope from the point to the radius what do i do next?

OpenStudy (amistre64):

we want an equation of the tangent line; so we have the slope of the tangent line; which point is the outside of the circle again?

OpenStudy (anonymous):

point Q(9,5) center (6,4)

OpenStudy (anonymous):

P(6,4)

OpenStudy (amistre64):

then all we have to do is fill in the exterior point in y = mx -mPx + Py y = -3x +3(9) + (5)

OpenStudy (amistre64):

the point slope form of a line; given (Px,Py) and a slope m is y - Py = m(x-Px) y = m(x-Px) + Py y = mx -mPx) + Py

OpenStudy (amistre64):

m=-3 in this case and we just fill in our given point (9,5)

OpenStudy (anonymous):

y-5=-3(x-9) ???